OpenStudy (ladiesman217):

help me please fans and medal awarded!!! What is the general form of the equation of a circle with center at (a, b) and radius of length m? x2 + y2 − 2ax − 2by + (a2 + b2 − m2) = 0 x2 + y2 + 2ax + 2by + (a2 + b2 − m2) = 0 x2 + y2 − 2ax − 2by + (a + b − m2) = 0 x2 + y2 + 2ax + 2by + a2 + b2 = -m2

OpenStudy (ladiesman217):

@mathslover can you help me please

mathslover (mathslover):

This is the very basic topic of "Circle" . I think, you should have a look at the lesson : http://www.mathsisfun.com/algebra/circle-equations.html

OpenStudy (ladiesman217):

i still dont understand it

jigglypuff314 (jigglypuff314):

the standard equation for a circle is: (x-a)^2 + (y-b)^2 = 0 then factor that and arrange, that will give you your answer

jigglypuff314 (jigglypuff314):

sorry typo: (x-a)^2 + (y-b)^2 = m^2 not 0

OpenStudy (ladiesman217):

i dont understand how to factor it thoug ive tried numerous times

jigglypuff314 (jigglypuff314):

(x-a)^2 + (y-b)^2 = m^2 means (x-a)(x-a) + (y-b)(y-b) = m^2 then you foil

OpenStudy (ladiesman217):

whatss foil

OpenStudy (ladiesman217):

ive only heard of pemdas

jigglypuff314 (jigglypuff314):

FOIL = First, Outer, Inner, Last for (x-a)(x-a) = first x*x outer x*(-a) inner (-a)*x last (-a)*(-a) so it equals x*x + x*(-a) + (-a)*x + (-a)(-a) so x^2 - 2ax + a^2

OpenStudy (ladiesman217):

ok thank you